40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16866     Entry time: Thu May 19 19:05:59 2022     In reply to: 16863     Reply to this: 16868
Author: Anchal 
Type: Update 
Category: BHD 
Subject: BS Chamber all work finished, BHD path setup 

[Anchal, Paco, Yuta]

BS Oplev Path

  • The changed position of PRM (40m/16863) meant that BS oplev path is getting clipped by the PRM SOS tower.
  • We had to move BSOL ~ 16 cm North and ~ 1.7 cm East.
  • This means that the BS Oplev input beam is now coming behind TT2 instead of infront of it.
  • We also had to align the beam such that input and returning beam are colinear.
  • This meant we, had to change the mount of the upstream beamsplitter in the in-air table so that we can use that for separating the return beam.
  • Again, we should order 2 inc visible BS for this path.
  • Half of the return beam is making its way all the way back to the laser head. I'm not sure if that can be an issue for our oplev loops.
  • We kept the SRM Oplev path same using irises on the table.

PRM Oplev

  • Again, due to changed position of PRM and BS Oplev, it became very hard to setup oplev for PRM.
  • We found a special position which allows us to catch returning beam through the center of the window.
    • But this returning beam is not prompt reflection from PRM, it is reflection of the HR surface.
    • We are hitting about ~5 mm from the edge of PRMOL mirror (because we cannot move the mirror anymore south to avoid clipping BS and SRM input oplev beams)
    • We put in a 1.1m focal length lens in the input beam to narrow the beam on PRMOL so that it doesn't clip
  • We did not put any lens for the return beam. The sensitivity of this oplev might be low due to slighlty bigger beam on the QPD than others (SRM, BS). We can revisit and insert a lens later if required.

Interferometer alignment and PRM alignment

  • The work on BS table did not change the table balance much. We got back the alignment pretty much instantly.
  • We were able to maximize the arm transmissions.
  • Then we used a beam card with hole to check for reflection from PRM and used PRM (mostly pitch correction) to get the return beam back in same way.
  • This recovered REFL beam on the camera. We used REFLDC signal to align PRM better and maximized it.
  • We centered BS, SRM and PRM oplevs after this point.

LO beam mode correction and spatial overlap

  • We tried changing the distance between LO3 and LO4 to get a better output LO beam.
  • We also tried to swap the LO4 mirror with the spare mirror but we had the same result.
  • Eventually, we decided to move LO3 back to East and LO4 to the west edge of the table. This made the beam sizes comparable.
    • Future exercise: We think that LO1 or LO2 might be significantly off-spec in their ROCs which might cause this issue.
    • We should rerun the calculations with the ROC values of LO1 and LO2 written in the datasheets and figure out the correct LO3-LO4 length required.
    • We can make this change in the next vent if required.
  • After the beam sizes were looking approximately similar but more iterations of changing length and realigning are required.

Remaining tasks before pumpdown

  • Push/pull too bright/dark OSEMs in the SOSs (40m/16865).
  • Finish LO beam mode correction and spatial overlap.
  • Center all oplevs, note all beam positions on camera, and note down all DC PD values at proper alignment.
ELOG V3.1.3-